The lines joining the origin to the points of intersection of the line lx + my = 1 and y2= 4ax are at right angles. Show that 4al = 1.


1
Expert's answer
2022-01-31T13:15:35-0500

The points of intersection of "lx+my=1" and "y^2=4ax" are:

"lx+my=1\Rightarrow my=1-lx\Rightarrow y=\dfrac{1-lx}m"

"\left( \dfrac{1-lx}m\right)^2=4ax\Rightarrow \dfrac{1-2lx+l^2x^2}{m^2}=4ax\Rightarrow"

"l^2x^2+(-2l-4am^2)x+1=0\Rightarrow x=\dfrac{l+2am^2\pm 2m\sqrt{a(l+am^2)}}{l^2}"

"\Rightarrow y=-2am\mp 2\sqrt{a(l+am^2)}"

The points of intersection are:

"P\left( \dfrac{l+2am^2+2m\sqrt{a(l+am^2)}}{l^2}, -2am-2\sqrt{a(l+am^2)}\right)=\overrightarrow{OP}"

"Q\left( \dfrac{l+2am^2-2m\sqrt{a(l+am^2)}}{l^2}, -2am+2\sqrt{a(l+am^2)}\right)=\overrightarrow{OQ}"

We know that the lines joining the origin "O" are at right angles: "\overrightarrow{OP}\cdot \overrightarrow{OQ}=0:"

"\dfrac{(l+2am^2+2m\sqrt{a(l+am^2)})(l+2am^2-2m\sqrt{a(l+am^2)})}{l^2\cdot l^2}+"

"+(-2am-2\sqrt{a(l+am^2)})(-2am+2\sqrt{a(l+am^2)})="

"=\dfrac{1}{l^2}-4la=0 \Rightarrow \dfrac{1}{l^2}=4la \Rightarrow 1=4lal^2\Rightarrow \boxed{1=4l^3a}"

We see that the problem statement is incorrect. It is correct that "4al^3=1" is satisfied.